OpenStudy (anonymous):

why cos^2(x)+sin^2(x)=1?

hero (hero):

Let Khan Academy explain it to you. Go there

OpenStudy (anonymous):

please give the link

OpenStudy (anonymous):

 ABC is a right triangle AC^2=AB^2+BC^2(pythagoras theorem) divide each term by AC^2 \[\frac{AB^2}{AC^2}+\frac{BC^2}{AC^2}=\frac{AC^2}{AC^2}\] \[(\frac{AB}{AC})^2+(\frac{BC}{AC})^2=1\] \[[\frac{AB}{AC}=\frac{adjacent}{hypotenuse}=cosx,\frac{BC}{AC}=\frac{opposite}{hypotenuse}=sinx]\] \[\therefore \cos^2x+\sin^2x=1\]

OpenStudy (anonymous):

Thank You

hero (hero):

Yeah but that still isn't a good explanation. How do you amazingly avoid explaining it in terms of the Unit Circle? I'm convinced this student has no clue what sine and cosine is in terms of the Unit Circle.